Tom sold a watch to Harry with a profit of 20% and Harry sold to Peter with a profit of 30%. If Peter paid $780, how much did Tom pay for the watch?
Answer by Cromlix(4381)

Hi there,
130% = $780
1% = 780/130
100% = 780/130 x 100
= $600
This is what Harry paid for the watch.
120% = $600
1% = 600/120
100% = 600/120 x 100
= $500
This is what Tom paid for the watch.
Hope this helps :-)